# The optimum mix of storage and backup in a highly renewable, highly reliable European electricity grid

### v2

@@ Abstract
We estimate the variability of solar and wind energy generation potential in Europe over a 43 year period between 1980–2022 with the MERRA-2 reanalysis datasets. We compare the estimated supply potential to hourly demand data from 36 European countries to calculate the reliability of a highly renewable electricity grid in Europe. We find that in cost-optimised scenarios with onshore wind, solar and storage, but no natural gas, reliably meeting the last 1% of demand represents 36% of the entire system cost. Including small amounts of dispatchable natural gas drastically reduces the cost of a renewable, highly reliable grid: overall system costs fall by 31% when just 1% of total generation is permitted to come from natural gas. Large renewable overbuild factors (greater than $\times$4 peak demand) are required to meet modern grid reliability standards in all scenarios, and wind, rather than solar, dominates the generation mix.
@@
Both storage and dispatchable backup are crucial for maintaining grid reliability
@@ 1 Introduction
Europe has some of the most aggressive CO2-reduction targets in the world, with the EU pursuing net zero carbon emissions by 2050 across its 27 member states [1] and several individual countries setting legally binding targets to reach the same milestone [2, 3, 4].
It is expected that most of the decarbonisation of the electricity sector will be achieved using wind and solar energy. To date, wind and solar have been incorporated into the European electricity grid with little effect on reliability, but maintaining modern reliability standards of 99.97% (meaning that demand exceeds supply for fewer than 3 hours per year on average) will be a major challenge for system operators as the share of weather-dependent generation grows [5].
@@
While no country-scale or continental-scale grid has ever operated with such high penetrations of wind and solar energy, there have been recent attempts to investigate the feasibility of a 100% renewable grid in Europe. Two prominent 2019 studies reported annual 2050 system costs of €276 billion [8] and €530 billion [9] respectively for a 100% renewable grid. Such studies generally estimate wind and solar generation potential across the European continent using climate reanalysis datasets, and feed this into an integrated power system model to optimise for the minimum-cost system that meets demand given constraints on generator capacity, transmission capacity, generator ramp-time and other factors. These models are computationally expensive and, as a result, are often limited in their temporal scope. The PLEXOS simulation in [9] cycled through one week of weather data per month for a single year, [^1] In the PLEXOS study [9], the authors did attempt to ensure that the year they analysed was the ‘worst’ weather year between 1979–2015 in terms of renewable generation. while the LUT model used in [8] also relied on a single year of weather data and assumed perfect foresight of supply and demand [10].
These steps are necessary to make the simulations tractable, but preclude a detailed study of grid reliability over multi-decadal time periods. Recent work by Ruggles et. al. has found that almost 40 years of weather data are required to ensure a reliable grid in systems without dispatchable backup [11]. Even when dispatchable generation is allowed to supply 5% of all electricity, 15 years of weather data are still required.
Computationally expensive models have other limitations. Previous studies have looked at the effects of carbon-intensity constraints and individual technology costs on the future grid [12, 13], but understanding how these factors combine to affect the overall system cost requires searching a high-dimensional space. For complex models, this can be prohibitively expensive.
In this paper, we give results from a reduced model for the European power sector capable of resolving some of these more difficult questions. Inspired by the work of Shaner et. al. [14] and Tong et. al. [15], we use the MERRA-2 reanalysis dataset to estimate solar and onshore-wind generation potential in Europe over a 43-year period from 1980–2022. We model Europe as a single node, averaging supply and demand across the entire continent and assuming perfect, lossless transmission. We then simulate this idealised grid with hourly resolution from 1980–2022, varying the installed capacities of solar, onshore wind, storage and dispatchable backup as exogenous inputs. With this single-node, four-technology grid model, we can investigate how three parameters—the permitted carbon intensity of the system, the cost of electricity storage, and the grid reliability target—interact to affect the overall system cost. These constraints are highly interdependent, and must be studied together in order to fully capture their impact on the system cost.
@@ 2 Methods I - Data
The simulations in this analysis rely on hourly time series data for solar generation and wind generation over a 43-year period (1980–2022). These time-series were generated using raw weather data from the MERRA-2 reanalysis dataset [16, 17]. Hourly electricity demand data for the year 2022 was collated from the ENTSO-E Transparency Platform [18].
@@ 2.3 Estimating electricity demand
The European Network of Transmission System Operators for Electricity (ENTSO-E), which operates the synchronous grid in Europe, requires its members to publish hourly electricity demand data every year, and makes this information publicly available [18]. However, not all countries connected to the synchronous European grid (CESA) are members of ENTSO-E, and the UK, Turkey and the North African states in particular do not publish hourly demand data on the ENTSO-E platform. Data for the UK are available elsewhere [27], but Turkey and the North African states were excluded from our analysis due to data being unavailable. Some countries, such as Ukraine and Albania, use the ENTSO-E platform but do not reliably publish data. We only included in our analysis the 36 countries that provided demand data for more than 8700 hours of the year (out of 8760) in 2022—see B. The borders of these countries define the locations of the MERRA-2 grid points used to calculate solar and wind generating potential.
@@ 2.4 Storage & Backup
3 Methods II - The Model
The result of the procedure detailed in Section 2 is a dataset containing average solar generation, average wind generation and average electricity demand at hourly resolution over the 43-year period. We use this dataset as an input to our simulation, which uses a rules-based dispatch model to calculate the average grid reliability and the percentage of electricity produced by natural gas for a given grid configuration. We then combine this with a cost model for system to find the minimum-cost blend given constraints on grid reliability and natural-gas usage (a proxy for carbon intensity if not paired with carbon capture and storage). We do not include transmission in our cost estimates, since we do not model transmission, but achieving the perfect levels of interconnectivity assumed here would require major upgrades to the European transmission grid.
@@ 3.2 System cost calculations
3.3 Limitations of Our Model
As discussed in the Introduction, questions of long-term grid reliability and of complex trade-offs between reliability, cost and carbon-intensity necessitate the use of simplified models. The single-node, four-technology model presented here allows these questions to be addressed, but it has some limitations that will now be discussed. Because we do not consider transmission and only model four technologies, the total system costs reported here are not realistic estimates for true European system costs. Specifically, including transmission would increase costs while including alternative and existing sources of generation would decrease the overall system cost.
@@ 5 Conclusions
We have simulated a renewables-dominated European electricity grid over a 43-year period, and have found that a reliable grid will require high levels of renewable overbuild—with installed peak generating capacity greater than four times the peak demand. Even so, we find that cost-minimised systems require both storage and dispatchable backup generation to achieve reliability targets. A large quantity of one cannot fully compensate for the complete absence of the other without incurring extreme penalties in system costs.

### v3

# Modelling reliability under deep decarbonisation of the European electricity grid

@@ Abstract
We introduce a new model able to characterise the relative costs of low-carbon generation mixes when subject to varying grid-reliability targets, which are evaluated using 43 years of historical weather data. The model can be used to develop cost-minimising strategies, and to assess the extent, duration, and frequency of outages under different technology mixes. For a 35-country subset of the European electricity grid, we show that reliably meeting the last 1% of demand accounts for about 36% of the entire system cost when only onshore wind, solar, and storage are used in the generation mix. We also show that by including small amounts of low-cost, dispatchable load-following generation (e.g., natural gas, hydro, or similar) the cost of a reliable, high-renewables grid can be drastically reduced. For example, the system cost can be reduced by 31% when just 1% of generation is allowed to come from natural gas.
@@
Finds that both storage and dispatchable backup are crucial for maintaining grid reliability
@@ 1 Introduction
Because of their low and falling costs, wind and solar power are expected to play a major role in the decarbonisation of electricity systems worldwide [1, 2, 3]. However, the inherent variability of wind and solar generation has led to questions about their potential to achieve very deep decarbonisation (near 100%) while also maintaining grid reliability at reasonable cost [4, 5, 6, 7, 8]. These concerns have motivated several studies on the feasibility of a 100% renewable power sector [9, 10, 11, 12, 13] and energy sector [3, 14, 15, 16] in Europe. Such studies generally estimate wind and solar generation potential across the European continent using climate reanalysis datasets, and feed this into a power system model to optimise for the minimum-cost system that meets demand given technical constraints on generator capacity, transmission capacity, generator ramp-time and other factors. However, these models are computationally expensive and, as a result, are often limited in their temporal scope. For example, two recent and widely-cited studies exploring 100% renewable scenarios in Europe employed the LUT energy system model [16] and the PLEXOS model energy system model [13]. Both studies were conducted using only a single year of weather data: the LUT study used a typical weather year [17], while the PLEXOS study used a performance-limiting weather year but included just one week of data per month during the year.
This reduction in temporal scope is necessary when using complex power sector models like PLEXOS and LUT, but it means that the results obtained are not robust over multi-decadal time scales. The need for high temporal resolution, long-duration weather data in energy system planning is already well-documented [18, 19, 20], and recent work by Ruggles et. al. has found that almost 40 years of weather data are required to ensure a reliable grid in systems without dispatchable backup [20]. Even when dispatchable generation is allowed to supply 5% of all electricity, 15 years of weather data are still required to ensure a reliable grid.
Computationally intensive models force other limitations as well. The LUT model assumes perfect foresight of supply and demand [17] in its simulations, which yields overly optimistic projections surrounding the levels of storage and backup needed to supplement a highly renewable grid. In addition, detailed sensitivity analyses are often impossible with computationally expensive models. While previous studies have explored the sensitivity of the system cost to single parameters like the permitted carbon intensity of the grid or individual technology costs [5, 21], scanning across multiple inputs simultaneously (e.g., investigating how varying the carbon intensity would affect the overall system cost under many different technology cost assumptions) becomes prohibitively expensive.
In order to overcome these challenges, we introduce here the RESCORE (Renewable Energy System Cost Optimisation and Reliability Evaluation) Model. Inspired by the work of Shaner et. al. [4] and Tong et. al. [22], RESCORE uses long-duration weather data to obtain accurate reliability estimates for highly-renewable grids. The model treats the grid as a single node, averaging supply and demand across the entire region of interest under the assumption of lossless transmission. RESCORE then simulates the single-node grid with hourly resolution from 1980–2022, varying the installed capacities of solar, onshore wind, storage, and dispatchable backup generation, and assuming a spatially homogenous distribution of renewable resources.
In this paper, we use RESCORE to give the first cost estimate of a highly-renewable grid in Europe with a realistic reliability constraint over a multi-decadal time period. Modern targets for grid reliability in Europe are 99.97% [23], but since desired grid reliability is an input to the RESCORE model, we are also able to explore for the first time how varying this reliability target would affect both the cost and make-up of a highly-renewable European grid. Furthermore, RESCORE returns full reliability information including the severity and duration of all outages, allowing these metrics to be analysed in detail. The speed of the model also allows for a rich sensitivity analyses in which the interplay between several different constraints and inputs can be investigated, which helps find global optimums faster. In this paper, we investigate the effect of simultaneously varying the target grid reliability, the carbon intensity, and the cost of electricity storage on the overall system cost. These constraints are highly interdependent, and must be studied together using a tool like RESCORE in order to fully capture their impact on the system cost.
@@ 2 Methods I - Data
The RESCORE model relies on hourly time series data for solar generation and wind generation over a 43-year period (1980–2022). These time-series were generated using raw weather data from the MERRA-2 reanalysis dataset [26, 27]. Hourly electricity demand data for the year 2022 was collated from the ENTSO-E Transparency Platform [28].
@@ 2.3 Estimating electricity demand
The European Network of Transmission System Operators for Electricity (ENTSO-E), which operates the synchronous grid in Europe, requires its members to publish hourly electricity demand data every year, and makes this information publicly available [28]. However, not all countries connected to the synchronous European grid (CESA) are members of ENTSO-E, and the UK, Turkey and the North African states in particular do not publish hourly demand data on the ENTSO-E platform. Data for the UK are available elsewhere [37], but Turkey and the North African states were excluded from our analysis due to data being unavailable. Some countries, such as Ukraine and Albania, use the ENTSO-E platform but do not reliably publish data. We only included in our analysis the 35 countries that provided demand data for more than 8700 hours of the year (out of 8760) in 2022—see B. The borders of these countries define the locations of the MERRA-2 grid points used to calculate solar and wind generating potential.
@@ 2.4 Storage & Backup
3 Methods II - The RESCORE Model
The result of the procedure detailed in Section 2 is a dataset containing average solar generation, average wind generation and average electricity demand at hourly resolution over the 43-year period. We use this dataset as an input to RESCORE, which uses a rules-based dispatch model to calculate the average grid reliability and the percentage of electricity produced by natural gas for a given grid configuration. We then combine this with a cost model for system to find the minimum-cost blend given constraints on grid reliability and natural-gas usage (a proxy for carbon intensity if not paired with carbon capture and storage). We do not include transmission in our cost estimates, since we do not model transmission, but achieving the perfect levels of interconnectivity assumed here would require major upgrades to the European transmission grid.
@@ 3.2 System cost calculations
3.3 Limitations of the RESCORE model
As discussed in the Introduction, questions of long-term grid reliability and of complex trade-offs between reliability, cost and carbon-intensity necessitate the use of simplified models. RESCORE allows these questions to be addressed, but it has some limitations that will now be discussed. Because we do not consider transmission and only model four technologies, the total system costs reported here are not realistic estimates for true European system costs. Specifically, including transmission would increase costs while including alternative and existing sources of generation would decrease the overall system cost.
@@ 5 Conclusions
We have simulated a renewables-dominated European electricity grid using the RESCORE model: new long-duration, high temporal resolution grid model over a 43-year period, and have found that a reliable grid will require high levels of renewable overbuild—with installed peak generating capacity greater than four times the peak demand. Even so, we find that cost-minimised systems require both storage and dispatchable backup generation to achieve reliability targets. A large quantity of one cannot fully compensate for the complete absence of the other without incurring extreme penalties in system costs.